Fermi National Laboratory
 

Accelerator Update for May 25, 2006 - May 26, 2006

 

Thursday May 25

The day shift began with Booster, Main Injector (MI), and the Tevatron (TeV) in access.

Operations reported at 11:50 AM that experts found the P3 line ground fault. Unfortunately, the fault was due to a bad magnet, a Main Ring B2 Dipole (HP3DS).

A Booster expert reported that their ORBMP work was finished. Operations established beam to Booster at 12:08 PM.

The Antiproton Source (Pbar) received permission to run beam at 4:03 PM.

Operations established beam to the MI dump at 4:57 PM. Experts began a study and tried to get beam to Pbar.

Operations established beam to the Recycler at 5:31 PM.

Operations reported at 5:56 PM that the TeV sector B11 separator was powered. Operators also reported that a hard ground fault was found between sectors A2 and A48.

Around 6:05 PM, operators accessed the TeV to search for the ground fault.

Pbar experts began a study at 8:02 PM.

TeV experts reported that the B11 separator had begun its bake.

The MI studies ended at 11:04 PM.

Friday May 26

The midnight shift began with Booster studies, with Pbar power supply problems, and with TeV power supply testing.

A power supply expert came in and repaired a Pbar shunt for HS925. Pbar resumed studies at 1:12 PM.

Operations reported at 2:17 AM that a MI magnet (MIQ611) had a high temperature alarm. An expert advised Operations to turn off the magnet and check the LCW valves. Operators accessed MI and at 3:31 AM found that the magnet was valved out.

Operations reestablished beam to MI at 4:08 AM.

The Plans for Today

The plans are to power the TeV this weekend and to get beam to Pbar. Beam commissioning should begin next week.

Machine Reports

Linac

Everything is running normally. The H- Source is providing a beam current of ~35 mA. The major down time during the week was due to LRF1 and KRF4 repairs. Experts tested a LINUX version of the Linac Steering Program. The collimators will be moved into the beam line next week.

Booster

The machine is running well. Experts have tuned across ORBMP, tuned the dump line, and worked on partial batching. They have found that the horizontal beam is too large; experts need some destructive beam tuning time to fix this problem. They will bump beam around the ring over the weekend and next week they will work on injection, the orbit, and transition.

MI

During the week experts commissioned power supplies, established beam at 8, 120, and 150 GeV, measured the MI-8 lattice, measured the 120 GeV lattice, moved MI quadrupole magnets to smooth the 120 GeV orbit, and established coalesced 150 GeV beam.

The plans for next week are to tune up for operations, send beam to users as requested, establish slip stacking, conduct aperture scans on the new large aperture quads, make MI lattice measurements, make harmonic corrections at 8 GeV, test LINUX programs, and possibly install BPMs and BLMs at MI-40.

Pbar

During this week, techs have rerouted PreVault sump water to the Main Ring, completed an OTR installation in AP1, started an A30 bake, began pulsing D:EKIK on Tuesday, conducted safety system tests, began running the power supplies on Wednesday, and conducted mock Recycler transfers. The target vault is ready for 120 GeV beam.

The priorities are as follows: measure the stacktail, establish reverse protons to the Debuncher, establish stacking beam, check Accumulator cooling with antiprotons, and to optimize the Debuncher acceptances and tune up stacking.

TeV

During this week, experts and techs have conditioned separators (2 new, 3 replaced), Hipotted the main bus, found a B11 separator vacuum leak, repaired a feed through, prepared to power main bus, low betas, and all correction circuits, and have tested the LINUX version of the Collider sequencer.

The Tevatron is cold.

Recycler

Problems with the Electron Cooling SF6 cooling skid has halted all electron tuning. Experts have stashed protons, and measured the lattice and lifetimes. Should be ready for antiprotons whenever Pbar is able to produce them.

Switchyard 120

Waiting for a B2 dipole magnet to be found to replace the bad one in the P3 line

MiniBooNE

Experts are ready for beam.

NuMI

The condensate tank pump has been repaired and tested. The target hall water mitigation was completed. The radiation monitoring installation was complete at EAV 1.5.

Work on the target chase chiller system commissioning and the target hall dehumidification system air balancing continues. There has been good progress on controls/instrumentation installations and on electrical installations. There will be test runs of the chiller and dehumidification systems thru the weekend.

Next Tuesday the following work will occur: the water and air sampling of NuMI will be completed, the re-core of NuMI, the search and secure of MI-65, and power supply testing.

There are low intensity beam and target scans scheduled for Wednesday.

Collider Detector at Fermilab

Experts reported that they had to pull the east plug to search for noisy channels. The work should be completed and the plug moved back in by next Wednesday, 5/31/06.

A wet engine failed. Cryo system techs should be working on that today.

D-Zero

Their detector check out continues with cosmic ray testing. Experts hope to power their magnet next Tuesday.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/26/2006    email Fermilab